Dog Up! Franks is looking at a new sausage system with an installed cost of $460,000. This cost will be depreciated straight-Please show me your clear work. Thank you so much!!
Answer #1

Cost of New Sausage system = $460,000

Project Life = 5 years

Annual Depreciation = Cost of New Sausage system/ Project Life

= 460,000/5

= $92,000

.

.

Profit on sale of Sausage system at the end of project = Sales Price – Book Value

= 59,000 – 0

= $59,000

After Tax Salvage Value of Sausage system at the end of project = Profit on sale of Sausage system at the end of project*(1-tax rate)

= 59.000*(1-0.21)

= 46,610

.

.

Calculation of NPV:

Year Inititial Investment (1) Working Capital (2) Savings in Pretax Operating Costs (3) After Tax Savings in Operating Cost (4) = (3)*(1-0.21) Depreciation (5) Tax Shield on Depreciation (6) = (5)*0.21 After tax Salvage Value (7) Net Cash Flows (8) = (1)+(2)+(4)+(6)+(7) DF Working Discounting Factor @9% (9) Present Value (10) = (9)*(8)
0                              (460,000)                         (26,500)                                                                          -                                                                                                              -                         -                                                                                   -                                                      -                                                                    (486,500)                      1                                                       1                                     (486,500.00)
1                                            -                                       -                                                                142,000                                                                                                112,180              92,000                                                                        19,320                                                    -                                                                      131,500 1/1.09^1 0.917431193                                        120,642.20
2                                            -                                       -                                                                142,000                                                                                                112,180              92,000                                                                        19,320                                                    -                                                                      131,500 1/1.09^2 0.841679993                                        110,680.92
3                                            -                                       -                                                                142,000                                                                                                112,180              92,000                                                                        19,320                                                    -                                                                      131,500 1/1.09^3 0.77218348                                        101,542.13
4                                            -                                       -                                                                142,000                                                                                                112,180              92,000                                                                        19,320                                                    -                                                                      131,500 1/1.09^4 0.708425211                                          93,157.92
5                                            -                             26,500                                                              142,000                                                                                                112,180              92,000                                                                        19,320                                           46,610                                                                    204,610 1/1.09^5 0.649931386                                        132,982.46
NPV:                                          72,505.62

.

NPV of this project is $72,505.62 . Since NPV is positive the project is viable.